# Exploring Holomorphic Retracts

### v1

@@ Conclusion
To address other hypotheses of the foregoing theorem briefly, let us remark that a non-trivial holomorphic retraction map is always a submersion except perhaps on a set of measure zero (cf. page $74$ of the book [Hrsch]). Nevertheless, we admit that it would be more significant to ascertain whether the origin falls outside such a measure zero set, for some retraction map $\rho$ as in the above theorem; or, if all retraction maps of such a class of domains is a submersion throughout the domain, making the submersive hypotheses redundant.
@@
Before going into the proof, we need to make some general remarks; to indicate as to why there are enough instances of domains verifying the conditions stipluated in proposition 1.6. First let us indicate as to why it may perhaps be of mild interest atleast, to note in passing that one may deduce out of the general facts about retracts and the already mentioned well-known characterization of retracts of polydiscs, yet another proof of the dramatic failure of the Riemann mapping theorem in multidimensional complex analysis namely, the biholomorphic inequivalence of two of the simplest bounded simply connected domains the ball $\mathbb{B}$ and $\Delta^{N}$ (despite their diffeomeorphic equivalence), as soon as the dimension $N$ gets any bigger than one; and other such inequivalences using retracts. We have already put this proof for a record in the ‘preliminaries’ section wherein we have formulated it more generally for a ball $D$ with respect to an arbitrary norm with $\mathbb{C}$-extremal boundary and product domains much more general than polydiscs (and so that proof is not repeated here). However here, we would indeed like to hasten to mention that this does not obtain for such $D$, results as good as those, which already well-known techniques in this realm can! – so, this leads to the question as to why one may want to even consider this proof even if the problem of biholomorphic inequivalence and its special case of establishing irreducibility of a domain $D$ i.e., showing that $D$ is not biholomorphic to a product of lower dimensional domains, is undoubtedly fundamental. To address this briefly here, let us recall that while the well-known techniques alluded to above, more explicitly such as that of Remmert – Stein (exposited well in [Encyclo-Vol6]; more specifically, see theorem $4.1$ therein alongwith the subsequent theorem $4.2$ due to Rischel) for showing irreducibility of domains, work for far more general domains $D$ beyond balanced domains of holomorphy, they impose the condition that $D$ not have any non-trivial analytic varieties in the boundary. We shall presently show that the route via retracts can very well dispense with such conditions, if $D$ is a balanced domain of holomorphy. To exposit this better by comparing and contrasting with the Remmert – Stein techniques, let us mention that the aforementioned proof by retracts for $\mathbb{B}\not\simeq\Delta^{N}$, shares (unlike a proof based on comparison of automorphism groups for instance), with the best possible techniques in this context, the essential idea about the difference in the lack or presence of analytic discs in the boundary of $D$ as being responsible for such inequivalences, as it has a bearing on the geometry of retracts through the origin in $D$. Moreover, this provides a shorter way to establish inequivalences without having to worry about the boundary extendibility (or boundary behaviour) of biholomorphic mappings since the boundary geometry is ‘internalized’ as features about the retracts through points in the interior; we believe this will be useful for establishing new inequivalences between pairs of bounded balanced domains in both of whose boundaries there reside non-trivial analytic discs. To illustrate this point with a concrete and simple example, consider the question of whether the analytic polyhedron
@@
when squared and combined with 7.1, leads to
@@
Let $D$ be any bounded domain in $\mathbb{C}^{N}$ with a Stein neighbourhood basis. Then any holomorphic retraction map on $D\setminus A$, where $A$ is a proper analytic subset of $D$ can be extended to a holomorphic retraction map on $D$. Consequently, every retract $R^{A}$ of $D\setminus A$ is of the form $R^{A}=R\cap(D\setminus A)$ for some retract $R$ of $D$.
@@
Firstly recall that removing an analytic set from a domain does not disconnect it; by definition $A$ is closed in $D$, thereby $D\setminus A$ is a domain in $\mathbb{C}^{N}$. Let $\rho:D\setminus A\to D\setminus A$ be a holomorphic retraction map. By applying the Riemann removable singularities theorem (for holomorphic functions of several complex variables) to each of the component functions of $\rho$ we get a (uniquely determined) holomorphic extension of $\rho$ which we denote by $\tilde{\rho}$. The fact that $\rho$ maps $D\setminus A$ into itself implies firstly that $\tilde{\rho}$ maps $D$ into its closure $\overline{D}$. However, it is easy to see that $\tilde{\rho}$ actually maps $D$ into itself. Indeed, pick any $a\in A$ and enquire if it is possible for $b=\tilde{\rho}(a)$ to lie in $\partial D$. Recalling the local structure of analytic sets, we know that: there exists a complex line segment $\ell\subset D$ through $a$ which intersects $A$ only at $a$. If $\varphi:\Delta\to D$ is a parametrization of $\ell$ with $\varphi(0)=a$, then $\tilde{\rho}\circ\varphi$ gives an analytic disc (with its image) intersecting $\partial D$ only at $D$. Since $D$ has a Stein neighbourhood basis, there exists an exhaustion function $\psi$ for $D$ which is plurisubharmonic on an open set containing $D$ in particular. Considering then the composition $\psi\circ\tilde{\rho}\circ\varphi$, we see that the maximum principle for this plurisubharmonic function is violated, under our assumption that $b\in\partial D$. Thus $\rho(D)$ lies within $D$ and as $a$ was arbitrary, we conclude that $\tilde{\rho}$ maps $A$ (and thereby $D$ as well) into $D$.
@@
We finish off by verifying that $\tilde{\rho}$ is a retraction on $D$. Indeed, we need only verify the required idempotence equation along $A$, as outside of $A$, $\tilde{\rho}$ is the given retraction map $\rho$ on $D\setminus A$. To do this, pick an arbitrary $a\in A$ and a sequence $(z_{n})\subset D\setminus A$ with $z_{n}$ converging to $a$. Then just by the idempotence of $\rho$ on $D\setminus A$, we have: $\rho\left(\rho(z_{n})\right)=\rho(z_{n})$, wherein by passing to limit using the continuity of $\tilde{\rho}$ twice, we get $\tilde{\rho}\left(\tilde{\rho}(a)\right)=\tilde{\rho}(a)$. As noted earlier, we may now conclude that the same equation holds not just on $A$ but throughout $D$ i.e., $\tilde{\rho}\circ\tilde{\rho}\equiv\tilde{\rho}$; denoting $\tilde{\rho}(D),\rho(D\setminus A)$ by $R^{A}$ and $R$ respectively, we arrive at the statement of the lemma. ∎
@@
We now detail the proof of proposition 1.5 and lay down some of its generalizations as well. This will be used in the sequel.
@@
Proof of proposition 1.5 . Let $p=(p_{1},p_{2})\in D_{1}\times D_{2}$ and $v_{2}\in T_{p_{2}}{D_{2}}$. Since $D_{1}$ is bounded, there exists $r_{1}>0$ such that $D_{1}\subset\Delta(p_{1},r_{1})$. Since $D_{2}$ is open, there exists $r_{2}>0$ such that $\Delta(p_{2},r_{2})\subset D_{2}$. Consider the domain $D:=D_{1}\times D_{2}$. Choose $v_{2}\in T_{p_{2}}D_{2}$ with $|v_{2}|<\frac{r_{2}}{r_{1}}$. Consider a holomorphic map $F:D_{1}\to\mathbb{C}$ defined by $F(z)=p_{2}+(z-p_{1})v_{2}$. Note that $F(p_{1})=p_{2}$ and $F^{\prime}(p_{1})=v_{2}$. Take $\epsilon=\frac{r_{2}}{r_{1}}$. If $|v_{2}|<\epsilon$, then $F$ maps $D_{1}$ into $D_{2}$. Consider a holomorphic retraction map $G:D_{1}\times D_{2}\to D_{1}\times D_{2}$ defined by $G(z,w)=(z,F(z))$. Hence $Z:=\{(z,F(z)):z\in D_{1}\}$ is a 1 dimensional retract of $D_{1}\times D_{2}$ passing through $(p_{1},p_{2})$ and $(1,v_{2})\in T_{p}Z$. ∎
@@
Suppose $D_{1},D_{2}$ be domains in $\mathbb{C}^{N},\mathbb{C}^{M}$ respectively and atleast one of domains is Caratheodory hyperbolic, say $D_{1}$; let $p=(p_{1},p_{2})$ be any given point in the product domain $D=D_{1}\times D_{2}$. Let $v_{1}\in T_{p_{1}}D_{1}$ be any non-zero vector. Then there exists $\epsilon>0$, such that for all vectors $v_{2}\in T_{p_{2}}D_{2}\cong\mathbb{C}^{M}$ with $\|v_{2}\|_{\infty}<\epsilon$, and there exists a retract $Z$ of $D_{1}\times D_{2}$ passing through $p$ such that $(w_{1},v_{2})\in T_{p}Z$, for some non-zero vector $w_{1}$ with $w_{1}\in span_{\mathbb{C}}\{v_{1}\}$.
@@
Since $D_{2}$ is open in $\mathbb{C}^{M}$, there exists a polydisk $\Delta^{M}(p_{2},r_{2})\subset D_{2}$. Since $D_{1}$ is Caratheodory hyperbolic, we have $\gamma_{D_{1}}(p_{1},v_{1})\neq 0$. This means that there exists a non-constant holomorphic map $g:D_{1}\to\Delta$ such that $g(p_{1})=0$ and $g^{\prime}(p_{1})(v_{1})\neq 0$. So $g^{\prime}(p_{1})$ is a non-zero linear map from $\mathbb{C}^{M}$ onto $\mathbb{C}$. Hence there exists $w_{1}\in T_{p_{1}}D_{1}$ with $w_{1}=\lambda v_{1}$ for some $\lambda\in\mathbb{C}$ and $g^{\prime}(p_{1})(w_{1})=1$. Choose $v_{2}\in T_{p_{2}}D_{2}$ with $v_{2}\in\Delta^{M}(0,r)$ Consider a holomorphic map $\phi:D_{1}\to\mathbb{C}^{M}$ defined by $\phi(z)=p_{2}+v_{2}g(z)$. Note that
@@
Write $r_{2}=(r_{2}^{(1)},\ldots,r_{2}^{(M)})$. Choose $\epsilon:=\min\{r_{2}^{(1)},\ldots,r_{2}^{(M)}\}$. If $\|v_{2}\|_{\infty}<\epsilon$, then $\phi$ maps $D_{1}$ into $D_{2}$. Note that $\phi(p_{1})=p_{2}$ and $\phi^{\prime}(p_{1})(w_{1})=v_{2}g^{\prime}(p_{1})(w_{1})=v_{2}$. Define a holomorphic map $F:D_{1}\times D_{2}\to D_{1}\times D_{2}$ by $F(z,w)=(z,\phi(z))$. Again note that $F(p_{1},w)=(p_{1},p_{2})$ and $F^{\prime}(p)(w_{1},w_{2})=(w_{1},\phi^{\prime}(p_{1})(w_{1}))=(w_{1},v_{2})$. Hence $Z:=\{(z,\phi(z)):z\in D_{1}\}$ is a retract of $D_{1}\times D_{2}$ passing through $(p_{1},p_{2})$ and $(w_{1},v_{2})\in T_{p}Z$, where $w_{1}\in span_{\mathbb{C}}<v_{1}>$. ∎
@@
With these observations, we are ready to prove the proposition by contradiction. Indeed, if $F$ were a biholomorphism mapping $D\setminus A$ onto a product domain $D$, then through $F(0)$ we have a retract along those directions given by proposition 7.2 – note that the hypothesis about Caratheodory hyperbolicity therein is verified owing to the fact that $D\setminus A$ is bounded; indeed thereby, every biholomorphic image of $D\setminus A$ must be Caratheodory hyperbolic as well. In particular, there always exists a non-trivial retract (necessarily of dimension one here) of $D$ for an open piece $S\subset\pi(S^{h})$ of directions at $F(0)$; here, $\pi$ denotes the projection onto the Euclidean sphere $\partial\mathbb{B}$ given by $\pi(z)=z/|z|_{l^{2}}$. By the biholomorphicity of $F$, it follows that for an open piece $S^{\prime}$ of directions, there always exists a (non-trivial) retract of $D$ through the origin with its tangent direction at 00 belonging to $S^{\prime}$. But this contradicts the hypotheses that $D$ is not convex at an open dense subset of $\partial D$ which implies that for each non-convex boundary point $a$ of $D$, there are no retracts of $D\setminus A$ passing through the origin in the direction of $a$. This finishes the proof. ∎

### v6

@@ Conclusion
To address other hypotheses of the foregoing theorem briefly, let us remark that a non-trivial holomorphic retraction map is always a submersion except perhaps on a set of measure zero (page $74$ of the book [Hrsch]). Nevertheless, we admit that it would be more significant to ascertain whether the origin falls outside such a measure zero set, for some retraction map $\rho$ as in the above theorem; or, if all retraction maps of such a class of domains is a submersion throughout the domain, making the submersive hypotheses redundant.
@@
Before we begin to work out the proof, we need to make some general remarks; to indicate as to why there are enough instances of domains verifying the conditions stipulated in proposition 1.6. Indeed, even before this, let us indicate as to why it may perhaps be of mild interest atleast, to note in passing (as we did in the preliminaries section) that: one may deduce out of the general facts about retracts and the already mentioned well-known characterization of retracts of polydiscs, yet another proof of the dramatic failure of the Riemann mapping theorem in multidimensional complex analysis which is, the biholomorphic inequivalence of two of the simplest bounded simply connected domains namely, the ball $\mathbb{B}$ and $\Delta^{N}$ (despite their diffeomeorphic equivalence), as soon as the dimension $N$ gets any bigger than one; and other such inequivalences using retracts. We have already put this proof for a record in the ‘preliminaries’ section wherein we have formulated it more generally for a ball $D$ with respect to an arbitrary norm with $\mathbb{C}$-extremal boundary and product domains much more general than polydiscs (and so that proof is not repeated here). However here, we would indeed like to hasten to mention that this does not obtain for such $D$, results as good as those, which already well-known techniques in this realm can! – so, this leads to the question as to why one may want to even consider this proof even if the problem of biholomorphic inequivalence and its special case of establishing irreducibility of a domain $D$ i.e., showing that $D$ is not biholomorphic to a product of lower dimensional domains, is undoubtedly fundamental. To address this briefly here, let us recall that while the well-known techniques alluded to above, more explicitly such as that of Remmert – Stein (exposited well in [Encyclo-Vol6]; more specifically, see theorem $4.1$ therein alongwith the subsequent theorem $4.2$ due to Rischel) for showing irreducibility of domains, work for far more general domains $D$ beyond balanced domains of holomorphy, they impose the condition that $D$ not have any non-trivial analytic varieties in the boundary. We shall presently show that the route via retracts can very well dispense with such conditions, if $D$ is a balanced domain of holomorphy. To exposit this better by comparing and contrasting with the Remmert – Stein techniques, let us mention that the aforementioned proof by retracts for $\mathbb{B}\not\simeq\Delta^{N}$, shares (unlike a proof based on comparison of automorphism groups for instance), with the best possible techniques in this context, the essential idea about the difference in the lack or presence of analytic discs in the boundary of $D$ as being responsible for such inequivalences, as it has a bearing on the geometry of retracts through the origin in $D$. Moreover, this provides a shorter way to establish inequivalences without having to worry about the boundary extendibility (or boundary behaviour) of biholomorphic mappings since the boundary geometry is ‘internalized’ as features about the retracts through points in the interior; we believe this will be useful for establishing new inequivalences between pairs of bounded balanced domains in both of whose boundaries there reside non-trivial analytic discs. To illustrate this point with a concrete and simple example, consider the question of whether the analytic polyhedron
@@
when squared and combined with 8.1, leads to
@@
Let $D$ be any bounded pseudoconvex domain in $\mathbb{C}^{N}$ either with a Stein neighbourhood basis or with a bounded plurisubharmonic defining function. Then any holomorphic retraction map on $D\setminus A$, where $A$ is a proper analytic subset of $D$ can be extended to a holomorphic retraction map on $D$. Consequently, every retract $R^{A}$ of $D\setminus A$ is of the form $R^{A}=R\cap(D\setminus A)$ for some retract $R$ of $D$.
@@
Firstly recall that removing an analytic set from a domain does not disconnect it; by definition $A$ is closed in $D$, thereby $D\setminus A$ is a domain in $\mathbb{C}^{N}$. Let $\rho:D\setminus A\to D\setminus A$ be a holomorphic retraction map. By applying the Riemann removable singularities theorem (for holomorphic functions of several complex variables) to each of the component functions of $\rho$ we get a (uniquely determined) holomorphic extension of $\rho$ which we denote by $\tilde{\rho}$. The fact that $\rho$ maps $D\setminus A$ into itself implies firstly that $\tilde{\rho}$ maps $D$ into its closure $\overline{D}$. However, it is easy to see that $\tilde{\rho}$ actually maps $D$ into itself. Indeed, pick any $a\in A$ and enquire if it is possible for $b=\tilde{\rho}(a)$ to lie in $\partial D$. Recalling the local structure of analytic sets, we know that: there exist a complex line segment $\ell\subset D$ through $a$ which intersects $A$ only at $a$. If $\varphi:\Delta\to D$ is a parametrization of $\ell$ with $\varphi(0)=a$, then $\tilde{\rho}\circ\varphi$ gives an analytic disc (with its image) intersecting $\partial D$ only at $a$. If $D$ admits a bounded plurisubharmonic defining function $\psi$, then we consider the composition $\phi:=\psi\circ\tilde{\rho}\circ\varphi$, which is a subharmonic function on $\Delta^{*}=\Delta\setminus\{0\}$; as it is also bounded (by the boundedness of $\psi$), it follows by that the origin is a removable singularity for this subharmonic function, thereby $\phi$ extends to a subharmonic function on $\Delta$ which we continue to denote by $\phi$. But then, the maximum principle for this subharmonic function is violated, under our assumption that $b\in\partial D$. We deduce therefore that $b$ cannot lie in $\partial D$ and it follows that $\tilde{\rho}(D)\subset D$. Even if the existence of such a $\psi$ is not known, it is indeed true that $\tilde{\rho}(D)$ lies within $D$ (the above argument argument was given to show a simple route under the aforementioned assumption about bounded plurisubharmonic defining functions). This follows from the last result in [Abe], owing to our hypothesis that $D$ has a Stein neighbourhood basis; as the work [Abe] has some non-major errors, we also cite the Masters thesis of Persson [Persson], wherein the errors are set right with a better exposition. Thus as $\tilde{\rho}(a)$ lies within $D$ and as $a$ was arbitrary, we conclude that $\tilde{\rho}$ maps $A$ (and thereby $D$ as well) into $D$.
@@
We finish off by verifying that $\tilde{\rho}$ is a retraction on $D$. Indeed, we need only verify the required idempotence equation along $A$ (as outside of $A$: $\tilde{\rho}$ is the given retraction map $\rho$ on $D\setminus A$). To do this, pick an arbitrary $a\in A$ and a sequence $(z_{n})\subset D\setminus A$ with $z_{n}$ converging to $a$. Then just by the idempotence of $\rho$ on $D\setminus A$, we have: $\rho\left(\rho(z_{n})\right)=\rho(z_{n})$, wherein by passing to limit using the continuity of $\tilde{\rho}$ twice, we get $\tilde{\rho}\left(\tilde{\rho}(a)\right)=\tilde{\rho}(a)$. As noted earlier, we may now conclude that the same equation holds not just on $A$ but throughout $D$ i.e., $\tilde{\rho}\circ\tilde{\rho}\equiv\tilde{\rho}$; denoting $\tilde{\rho}(D),\rho(D\setminus A)$ by $R^{A}$ and $R$ respectively, we arrive at the statement of the lemma. ∎
Remark 8.2.
One may want to know to what extent we may weaken the hypotheses of the above lemma. We do not pursue this here but draw attention to the fact that rather simple counterexamples exist to show that boundedness cannot be dropped easily. Indeed, consider the unbounded domain $D=\mathbb{C}^{2}\setminus A$, where $A=\{z=0\}$, the $w$-axis. Notice then that $R^{A}=\{(z,1/z):z\in\mathbb{C}^{*}\}$ is a retract of $D$. We observe that there exists no retract $R$ of $\mathbb{C}^{2}$ such that $R^{A}=R\cap(D\setminus A)$. Indeed, this follows from the fact that the (smooth) analytic variety $R^{A}$ described by the equation $zw-1=0$, is irreducible, together with the principle of analytic continuation for complex analytic sets.
Notwithstanding the example in the above remark, we may certainly remove analytic varieties of higher codimension to attain the analogue of the foregoing lemma for the case when the domain is the entire complex Euclidean space itself: to adapt the essential ideas of its proof to deal with the unboundedness now, we only need to employ (theorem 5B in [Whitney]) in place of the standard Riemann removable singularities theorem. We shall not repeat the details and be content with stating it as follows.
Corollary 8.3.
(Corollary to the proof of lemma 8.1) Let $A$ be any analytic variety of codimension at-least two in $\mathbb{C}^{N}$. Then every retract $R^{A}$ of $\mathbb{C}^{N}\setminus A$ is of the form $R^{A}=R\cap(\mathbb{C}^{N}\setminus A)$ for some retract $R$ of $\mathbb{C}^{N}$.
@@
Without further ado now, we proceed to detailing the proof of proposition 1.5 and lay down some of its generalizations as well. This will be used in the sequel.
@@
Proof of proposition 1.5 . Let $p=(p_{1},p_{2})\in D_{1}\times D_{2}$ Since $D_{1}$ is bounded, there exist $r_{1}>0$ such that $D_{1}\subset\Delta(p_{1},r_{1})$. Since $D_{2}$ is open, there exist $r_{2}>0$ such that $\Delta(p_{2},r_{2})\subset D_{2}$. Consider the domain $D:=D_{1}\times D_{2}$. Choose $v_{2}\in T_{p_{2}}D_{2}$ with $|v_{2}|<\frac{r_{2}}{r_{1}}$. Consider a holomorphic map $F:D_{1}\to\mathbb{C}$ defined by $F(z)=p_{2}+(z-p_{1})v_{2}$. Note that $F(p_{1})=p_{2}$ and $F^{\prime}(p_{1})=v_{2}$. Take $\epsilon=\frac{r_{2}}{r_{1}}$. If $|v_{2}|<\epsilon$, then $F$ maps $D_{1}$ into $D_{2}$. Consider a holomorphic retraction map $G:D_{1}\times D_{2}\to D_{1}\times D_{2}$ defined by $G(z,w)=(z,F(z))$. Hence $Z:=\{(z,F(z)):z\in D_{1}\}$ is a one-dimensional retract of $D_{1}\times D_{2}$ passing through $(p_{1},p_{2})$ and $(1,v_{2})\in T_{p}Z$. ∎
@@
Suppose $D_{1},D_{2}$ be domains in $\mathbb{C}^{N},\mathbb{C}^{M}$ respectively and atleast one of domains is Caratheodory hyperbolic, say $D_{1}$; let $p=(p_{1},p_{2})$ be any given point in the product domain $D=D_{1}\times D_{2}$. Let $v_{1}\in T_{p_{1}}D_{1}$ be any non-zero vector. Then there exist $\epsilon>0$, such that for all vectors $v_{2}\in T_{p_{2}}D_{2}\cong\mathbb{C}^{M}$ with $\|v_{2}\|_{\infty}<\epsilon$, there is a retract $Z$ of $D_{1}\times D_{2}$ passing through $p$ such that $(w_{1},v_{2})\in T_{p}Z$, for some non-zero vector $w_{1}$ with $w_{1}\in span_{\mathbb{C}}\{v_{1}\}$.
@@
Since $D_{2}$ is open in $\mathbb{C}^{M}$, there exist a polydisk $\Delta^{M}(p_{2},r_{2})\subset D_{2}$. Since $D_{1}$ is Caratheodory hyperbolic, we have $\gamma_{D_{1}}(p_{1},v_{1})\neq 0$. This means that there exist a non-constant holomorphic map $g:D_{1}\to\Delta$ such that $g(p_{1})=0$ and $g^{\prime}(p_{1})(v_{1})\neq 0$. So $g^{\prime}(p_{1})$ is a non-zero linear map from $\mathbb{C}^{M}$ onto $\mathbb{C}$. Hence there exist $w_{1}\in T_{p_{1}}D_{1}$ with $w_{1}=\lambda v_{1}$ for some $\lambda\in\mathbb{C}$ and $g^{\prime}(p_{1})(w_{1})=1$. For $v_{2}\in T_{p_{2}}D_{2}$, consider the holomorphic map $\phi:D_{1}\to\mathbb{C}^{M}$ defined by $\phi(z)=p_{2}+v_{2}g(z)$. Note that
@@
Write $r_{2}=(r_{2}^{(1)},\ldots,r_{2}^{(M)})$. Let $\epsilon:=\min\{r_{2}^{(1)},\ldots,r_{2}^{(M)}\}$. Choose $v_{2}\in T_{p_{2}}D_{2}$ with $\|v_{2}\|_{\infty}<\epsilon$. Hence $\phi$ maps $D_{1}$ into $D_{2}$. Note that $\phi(p_{1})=p_{2}$ and $\phi^{\prime}(p_{1})(w_{1})=v_{2}g^{\prime}(p_{1})(w_{1})=v_{2}$. Consider the holomorphic map $F:D_{1}\times D_{2}\to D_{1}\times D_{2}$ defined by $F(z,w)=(z,\phi(z))$. Again note that $F(p_{1},w)=(p_{1},p_{2})$ and $F^{\prime}(p)(w_{1},w_{2})=(w_{1},\phi^{\prime}(p_{1})(w_{1}))=(w_{1},v_{2})$. Hence $Z:=\{(z,\phi(z)):z\in D_{1}\}$ is a retract of $D_{1}\times D_{2}$ passing through $(p_{1},p_{2})$ and $(w_{1},v_{2})\in T_{p}Z$, where $w_{1}\in span_{\mathbb{C}}<v_{1}>$. ∎
@@
With these observations, we are ready to prove the proposition by contradiction. Indeed, if $F$ were a biholomorphism mapping $D\setminus A$ onto a product domain $D$, then through $F(0)$ we have a retract along those directions given by proposition 8.4 – note that the hypothesis about Caratheodory hyperbolicity therein is verified owing to the fact that $D\setminus A$ is bounded; indeed thereby, every biholomorphic image of $D\setminus A$ must be Caratheodory hyperbolic as well. In particular, there always exist a non-trivial retract (necessarily of dimension one here) of $D$ for an open piece $S\subset\pi(S^{h})$ of directions at $F(0)$; here, $\pi$ denotes the projection onto the Euclidean sphere $\partial\mathbb{B}$ given by $\pi(z)=z/|z|_{l^{2}}$. By the biholomorphicity of $F$, it follows that for an open piece $S^{\prime}$ of directions, there always exist a (non-trivial) retract of $D$ through the origin with its tangent direction at 00 belonging to $S^{\prime}$. But this contradicts the hypotheses that $D$ is not convex at an open dense subset of $\partial D$ which implies that for each non-convex boundary point $a$ of $D$, there are no retracts of $D\setminus A$ passing through the origin in the direction of $a$. This finishes the proof. ∎